the hight and distance along horizontal for body projected in vertical plane are given by y=8t-5t2 and x=6t so what is it initial velocity?

the hightand distancealong horizontal for body projected in vertical plane are given by y=8t-5t2 and x=6t so what is it initial velocity?

Grade:11

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Nivedita

for calculation of velocity differentiate with time and then put t=0 for initial velocity.

Vy= 8 - 10t = 8

Vx= 6

hence V= √82 + 62  = 10  Answer

All the best

AKASH GOYAL
AskiitiansExpert-IITD

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.